OpenStudy (anonymous):

antiderivative of negative 2 over x cubed

OpenStudy (anonymous):

\[\int -\frac{2}{x^3}\] You can rewrite that as \[-2\int x^{-3}\] Which should hopefully help make it a lot simpler. :D